# Green Federated Learning: A new era of Green Aware AI

### v1

@@ Abstract.
Artificial intelligence, Green federated learning
ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Green federated learning
1. Introduction
1.1. Contest Scenario
@@
Since 2012, the discipline of artificial intelligence of artificial intelligence (AI) has made impressive strides in various areas, including speech recognition, object recognition, machine translation, and game playing. Machine learning models that are getting bigger and more computationally intensive have contributed significantly to this advancement. Initially, AI research focuses on the accuracy improvement of the model using enormous computational energy, leading to Red AI (Schwartz et al., 2020). Hence, the computational cost is very high, and due to massive power consumption, Red AI is not environment-friendly, leading to “Green AI” - AI research that is more inclusive and friendly to the environment. The phrase “Green AI” relates to AI research that produces new insights while accounting for the computational cost, promoting decreased resource use. In contrast to Red AI, Green AI encourages methods with favorable performance/efficiency trade-offs, which has led to fast-rising computational (and consequently carbon) costs. Researchers will have the choice to focus on the efficiency of their models with a positive influence on both inclusiveness and the environment if efficiency measurements are widely acknowledged as essential assessment metrics for research along with accuracy. Traditional machine learning approaches are unsuitable for green AI, as traditional ML algorithms need centralized data collection and processing by a centralized server, which is becoming a bottleneck of large-scale applications in daily life due to progressively growing privacy issues.
@@
“Federated learning” (FL) is a recent distributed learning paradigm where a central server coordinates a global model’s cooperative training by several clients. Federated learning keeps the training data local while allowing resource-constrained edge computing devices, like IoT and smartphones, to create a common model for prediction. Clients must submit updated models (such as those represented by gradients or parameters) to the server after training the models on their private local data. Cross-device FL and cross-silo FL are two categories of FL that can be distinguished by the customers who participate and the training scale. Clients in cross-device FL are tiny, dispersed entities (smartphones, wearables, and edge devices, for example), and each client probably has a modest quantity of local data. Therefore, for cross-device FL to succeed, a sizable number of edge devices—up to millions—must participate in the training process. However, in cross-silo FL, clients are usually businesses or institutions (banks, hospitals, etc.). There are very few participants—between two and one hundred, perhaps—and every client is required to be present for the duration of the session (Huang et al., 2022).
@@
Federated Learning (FL) may also be resource-intensive and have a high energy consumption, especially when used on a scale. In the smart environment, billions of IoT devices, including smartphones, laptops, wearable technology, and cars, are connected by IoT networks (Al-Abiad et al., 2021). FL at the edge should concentrate on reducing latency and resource energy usage without compromising the global model’s convergence rate in order to boost system reliability and advance green IoT. In terms of energy-efficient FL, reducing the energy consumption of edge devices is crucial to maintaining the long-term sustainability of the FL framework because mobile devices run on batteries. Reducing communication and computing energy should get significant consideration in a green FL or energy-efficient system. To reduce energy consumption and achieve an optimized model in terms of performance and efficiency, Green Federated Learning is introduced by the research community, which ultimately promotes sustainable AI. Green federated learning addresses two significant issues - (i) Privacy concerns: a central server, and local end devices maintain the same model by exchanging model updates instead of raw data, with which the privacy of data stored on end devices is not directly revealed, and (ii) Huge energy consumption issue: Green FL involves choosing design elements and FL parameters to decrease energy consumption while maintaining competitive performance and training time.
1.2. Motivation
@@
Focus Survey Heterogeneity Efficiency Application Green AI Statistical System Communication Computation A1 A2 A3 B1 B2 B3 P S Q (Li et al., 2020b) X X X X X X X X (Lim et al., 2020a) X X X X X X X X (Kairouz et al., 2021) X X X X X X X X X X (Wahab et al., 2021) X X X X X X X X X (Rahman et al., 2021) X X X X X X X X (Bellavista et al., 2021) X X X X X X X X (Gao et al., 2022) X X X X X X X X X X (Ye et al., 2023) X X X X X (Gecer and Garbinato, 2024) X X X X X X X X Ours
@@
Numerous surveys of federated learning in general or particular facets have been made available. Still, there is a need to offer a sensible and thorough taxonomy of the current state of energy-efficient (green) federated learning research difficulties and state-of-the-art, which is a significant and developing area of research for green-aware AI. Thus, we highlight the distinctive contributions of our work in Table 1 and evaluate the important contributions and limitations of previous relevant studies. The difficulties posed by federated learning are examined by Li et al.(Li et al., 2020b) from four angles: statistical heterogeneity, system heterogeneity, communication efficiency, and privacy concerns. They also provide a quick overview of a number of potential future study areas. In the study of federated learning in mobile edge networks, Lim et al.(Lim et al., 2020a) separated the approaches now in use into two categories: those that address the core issues with federated learning and those that apply federated learning to address edge computing issues. In addition to discussing current developments in federated learning, Kairouz et al.(Kairouz et al., 2021) offered an overview of unresolved issues and concerns, such as effective communication, privacy protection, attack defense, and federated fairness. Wahab et al.(Wahab et al., 2021) offer a detailed categorization system of current problems and solutions. The work mainly considered client selection and communication efficiency approaches. Rahaman et al.(Rahman et al., 2021) divided the survey’s categories into data partitioning, FL architectures, algorithms/aggregation techniques, and personalization techniques as the implementation details and FL applications in various industries and domains, and communication cost, statistical heterogeneity, systems heterogeneity, and privacy/security as the main challenges. In (Bellavista et al., 2021), the authors included information on the characteristics of federated environments, a unique taxonomy of decentralized learning methodologies, and a thorough explanation of the most pertinent and particular system-level contributions made by the surveyed solutions for non-IIDness, privacy, device heterogeneity, and poisoning defense. Data space, statistics, systems, and model heterogeneity in FL are covered by Gao et al.(Gao et al., 2022), who also offer a taxonomy and an overview of the scenarios, objectives, and techniques associated with each heterogeneity challenge. In addition to offering multiple viewpoints to show the potential for future development of heterogeneous federated learning, the authors of in (Ye et al., 2023) developed a new taxonomy to classify current heterogeneous federated learning approaches into three levels: data-level, model-level, and server-level. Gecer et al.(Gecer and Garbinato, 2024) conducted a survey of FL concerning efficiency, efficacy, privacy, robustness, and fairness. In order to compare and characterize federated learning solutions, the authors identified three abstract roles to present their conceptual model: aggregator, learner, and data creator. Some of the previous survey works (Ma et al., 2023; Pfeiffer et al., 2023; Qi et al., 2024) are related to specific concerns such as non-IID data, heterogeneity, model aggregation etc.
@@
The primary research question (RQ1) focuses on popularly used FL variants and the roadmap of green FL. The RQ1 is addressed by sections 3 and 5. The second research question (RQ2) is to offer a thorough and organized summary of all literature pertaining to green FL. Additionally, (RQ2) intends to demonstrate how adopting FL might be advantageous for the IoT application. The third research question (RQ3) is also driven by the need to address issues with energy efficiency and optimization in green FL IoT contexts. The RQ2 and RQ3 are addressed by 6. The fourth research question (RQ4) offers final recommendations for a researcher in the green FL area that are mostly focused on issues with IoT applications. RQ4 is addressed by the sections 7 and 8. In this survey work, we explore FL approaches and the problems of an energy-efficient optimized distributed model, as well as challenges and opportunities with the application of the Internet of Things (IoT). It is pertinent to mention that due to ample evidence of security and privacy-preserving challenges and solutions, in this survey, we are not considering the security and privacy-preserving issues.
1.3. Main Contributions
@@ 2. Research Method
The research methodology used for this study is the “preferred reporting item for systematic reviews and meta-analyses” (PRISMA) (Page et al., 2021). The PRISMA technique is a widely used standard for providing evidence in systematic reviews, and it has been embraced by journals and organizations relevant to health. PRISMA methodologies provide several benefits, including enhancing the quality of the review, allowing the reader to evaluate its merits and shortcomings, simulating review procedures, and structuring and formatting the review using PRISMA headings. However, conducting a comprehensive examination and releasing it can take some time. It can also become rapidly obsolete, so it needs to be updated regularly to include all new primary data released since the start of the project.
@@ 2.2. Selection Criteria
3. Challenges and solutions in Federated Learning
@@ 3.1. Optimization of Effectiveness
1) System heterogeneity: System parameters like hardware, network connectivity, and energy limitations frequently differ between different clients (Wei et al., 2022). The straggler effect may result from this, namely in cross-device FL.
2) Statistical heterogeneity is related to client non-i.i.d. data (Zhu et al., 2021). For instance, different hospitals could have non-identifiable disease data because their patient population varies in geography and demographics. Different banks may have heterogeneous data because clients from various areas and nations may have varied financial backgrounds and investing styles. Three key approaches have been created by current studies to address statistical heterogeneity, and they are presented as follows:
@@
Personalization: Due to the drift of heterogeneous data, the global model generated by the server may hurt the performance and generalization of the local models on the clients. To solve this problem, personalization techniques modify the local models according to the local tasks. Various techniques are available for personalization. The approaches to this kind of personalization are either fine-tuning a global model on a local client distribution or pruning a global model. Nevertheless, these current approaches either predetermine network layers for fine-tuning, resulting in inadequate global information storage inside client models, or personalize at the expense of critical global knowledge retention (Tamirisa et al., 2023). Another approach for personalization is personalization layers. The negative consequences of statistical heterogeneity can be mitigated using personalization layers not subject to the federated averaging (FedAvg) process (Arivazhagan et al., 2019a). Another approach for personalization is knowledge distillation, a technique for knowledge transfer, to compare the results of a shared dataset and learn from each other’s inferred knowledge (Jeong and Kountouris, 2023).
Clustering: Though most FL frameworks operate under the assumption that there is only one global model, learning clients’ knowledge may not be adequate in extremely heterogeneous data sets. Clustering techniques address this problem by dividing the clients into several clusters and ensuring that each cluster upholds a server-client architecture. Specifically, similar clients are assigned to the same cluster based on the loss values or submitted model weights that indicate how similar they are (Ouyang et al., 2021).
3.2. Optimization of Efficiency
@@ 3.2.1. Communication Overhead
The literature mentions two main approaches to reduce communication overhead in federated learning. They are model compression and communication rounds (Ángel Morell et al., 2022). Client selection is another approach to reduce communication overhead as the involvement of more client participation makes convergent the global model earlier. Reducing the number of global iterations by synchronizing the maximum number of clients eventually lowers communication expenses and energy usage. (Thakur et al., 2023a).
@@
Model compression can be achieved either using quantization or sparsification. Sparsification significantly lowers communication overhead, but it must be used carefully since transmitting too little information might negatively impact accuracy, while transmitting too much information can result in energy depletion and network congestion. Consequently, there is a trade-off between excessive and insufficient compression. It is costly to try to find the ideal balance through hyperparameter tweaking, and the tuning parameters that are produced are static and problem-specific. Furthermore, the majority of current compression techniques don’t take into account the varying communication costs associated with various technologies (Khirirat et al., 2021). Furthermore, because edge devices in the Internet of Things (IoT) have high energy constraints, the compression framework should be created to guarantee that the greatest amount of learning can occur while utilizing the least amount of energy to transmit the models—that is, to maximize the energy efficiency of edge nodes.
@@
Sparsification is a model compression technique for automatically determining the near-optimal communication and computation trade-off (Han et al., 2020). This trade-off is adjusted in the original FL method, federated averaging (FedAvg), by the number of local update rounds that occur in between each of the two communication (weight aggregation) rounds (McMahan et al., 2016). FedAvg either delivers all of the model parameters or nothing at all following each local model update step using gradient descent. Gradient sparsification (GS), a more balanced method that sends a sparse vector containing a selection of significant values from the complete gradient, has drawn interest recently in remote learning systems (Jiang and Agrawal, 2018). In contrast to FedAvg’s ”send-all-or-nothing” strategy, GS offers more flexibility in managing the trade-off between computation and communication (Han et al., 2020). However, these approaches employ a single compression rate for every user and fail to take into account the communication heterogeneity in an actual FL system; consequently, these approaches are constrained by the worst communication capacity among users. Moreover, sparsification approaches lack adaptability and fail to leverage the redundant, similar information in users’ ML models for compression. In (Beitollahi et al., 2022), the authors presented a novel approach to Dynamic Sparsification for Federated Learning (DSFL) that allows users to compress their local models according to their communication capacity at each iteration through the use of two novel methods for sparsification: layer-wise similarity sparsification (LSS) and extended top-K sparsification. DSFL is made possible by LSS to take advantage of the global redundant information present in users’ models through the sparsification process of Centralized Kernel Alignment (CKA) similarity. By allowing various values of the sparsification rate K for each user at each iteration, the extended top-K model sparsification approach enables DSFL to accommodate the diverse communication capability of user devices. Still, maintaining accuracy following a high ratio sparsification has never been addressed. To rectify the sparse gradient update process, a General Gradient Sparsification (GGS) framework is put forth in (Li et al., 2020a) for adaptive optimizers. Gradient correction and batch normalization update with local gradients (BN-LG) are its two main techniques. The optimizer can correctly handle the accumulated unimportant gradients with gradient correction, improving the model’s convergence. Moreover, local gradient updates to the batch normalization layer can mitigate the effects of delayed gradients without adding to the communication overhead.
@@
Quantization is another widely used model compression technique. The FL setup presents several difficulties. For instance, depending on the requirements for their hardware design, various clients may have varying computational limitations. The hardware accelerator’s supported quantization bit-width is one realistically significant heterogeneous hardware attribute. A well-trained model ought to be quantized to a range of bit widths that are representative of the heterogeneous device landscape without exhibiting appreciable performance reduction. The goal of quantization robustness is to do this by training a single model with the constraint of robustness to different quantization bit-widths while at inference time, without the need for retraining or fine-tuning (Gupta et al., 2023a). The research hasn’t given enough thought to the energy requirements and hardware-induced limits for on-device learning, despite the effectiveness of common Federated optimization techniques like Federated Averaging (FedAvg) in FL. Enabling trained models to be quantized to different bit-widths based on the energy requirements and heterogeneous hardware designs across the federation is specifically a requirement for on-device learning.
@@
Nevertheless, these approaches, such as sparsification and quantization are pre-defined. They fail to capture redundant, correlated information across FL iterations, user device data, and ML model parameters. Furthermore, the error-correcting potential of the FL process is not entirely utilized by these methods. In (Beitollahi and Lu, 2022), the authors presented the Federated Learning with Autoencoder Compression (FLAC) method, which compresses user devices’ models for uplink transmission by taking advantage of FL’s redundant information and error-correcting capabilities. In the Training State, FLAC trains an autoencoder on the server to encode and decode user models. Later, in the Compression State, the autoencoder is sent to user devices to compress local models for further iterations. FLAC dynamically manages the autoencoder error by alternating between the Training State and Compression State to modify its autoencoder and its compression rate based on the error tolerance of the FL system, ensuring the convergence of the FL.
@@
Client Selection is another method to address the communication efficiency challenge in FL. Transferring a large number of updated model parameters over the uplink communication channel from the users to the server is a challenging task in FL which increases the communication complexity is a challenging task. Reducing the number of participating clients by scheduling policies (Amiri et al., 2020; Yang et al., 2020) can help address this difficulty. In the literature, only a subset of clients can be scheduled for updates at each iteration due to bandwidth limitations. However, the random selection of a partial number of clients can reduce the convergence speed in the presence of heterogeneous clients. Moreover, selecting more clients ensures fast convergence in the presence of heterogeneous clients (Huang et al., 2023a). Hence, it is necessary to select participating clients to maintain the tradeoff between accuracy, convergence speed, and communication efficiency.
3.2.2. Computation Overhead
@@
4. Preliminaries and Problem Definition
@@
5. Variants of Federated Learning Frameworks

### v2

@@ Abstract.
footnotetext: DIMES - University of Calabria, Italy and Department of Mathematics and Applications ”R. Caccioppoli”, University of Naples Federico II, Italy
Introduction
Contest Scenario
@@ Introduction
Since 2012, the discipline of artificial intelligence of artificial intelligence (AI) has made impressive strides in various areas, including speech recognition, object recognition, machine translation, and game playing. Machine learning models that are getting bigger and more computationally intensive have contributed significantly to this advancement. Initially, AI research focuses on the accuracy improvement of the model using enormous computational energy, leading to Red AI [Schwartz et al.(2020)]. Hence, the computational cost is very high, and due to massive power consumption, Red AI is not environment-friendly, leading to “Green AI” - AI research that is more inclusive and friendly to the environment. The phrase “Green AI” relates to AI research that produces new insights while accounting for the computational cost, promoting decreased resource use. In contrast to Red AI, Green AI encourages methods with favorable performance/efficiency trade-offs, which has led to fast-rising computational (and consequently carbon) costs. Researchers will have the choice to focus on the efficiency of their models with a positive influence on both inclusiveness and the environment if efficiency measurements are widely acknowledged as essential assessment metrics for research along with accuracy. Traditional machine learning approaches are unsuitable for green AI, as traditional ML algorithms need centralized data collection and processing by a centralized server, which is becoming a bottleneck of large-scale applications in daily life due to progressively growing privacy issues.
@@
“Federated learning” (FL) is a recent distributed learning paradigm where a central server coordinates a global model’s cooperative training by several clients. Federated learning keeps the training data local while allowing resource-constrained edge computing devices, like IoT and smartphones, to create a common model for prediction. Clients must submit updated models (such as those represented by gradients or parameters) to the server after training the models on their private local data. Cross-device FL and cross-silo FL are two categories of FL that can be distinguished by the customers who participate and the training scale. Clients in cross-device FL are tiny, dispersed entities (smartphones, wearables, and edge devices, for example), and each client probably has a modest quantity of local data. Therefore, for cross-device FL to succeed, a sizable number of edge devices—up to millions—must participate in the training process. However, in cross-silo FL, clients are usually businesses or institutions (banks, hospitals, etc.). There are very few participants—between two and one hundred, perhaps—and every client is required to be present for the duration of the session [Huang et al.(2022)].
@@
Federated Learning (FL) may also be resource-intensive and have a high energy consumption, especially when used on a scale. In the smart environment, billions of IoT devices, including smartphones, laptops, wearable technology, and cars, are connected by IoT networks [Al-Abiad et al.(2021)]. FL at the edge should concentrate on reducing latency and resource energy usage without compromising the global model’s convergence rate in order to boost system reliability and advance green IoT. In terms of energy-efficient FL, reducing the energy consumption of edge devices is crucial to maintaining the long-term sustainability of the FL framework because mobile devices run on batteries. Reducing communication and computing energy should get significant consideration in a green FL or energy-efficient system. To reduce energy consumption and achieve an optimized model in terms of performance and efficiency, Green Federated Learning is introduced by the research community, which ultimately promotes sustainable AI. Green federated learning addresses two significant issues - (i) Privacy concerns: a central server, and local end devices maintain the same model by exchanging model updates instead of raw data, with which the privacy of data stored on end devices is not directly revealed, and (ii) Huge energy consumption issue: Green FL involves choosing design elements and FL parameters to decrease energy consumption while maintaining competitive performance and training time.
Motivation
@@
Focus Survey Heterogeneity Efficiency Application Green AI Statistical System Communication Computation A1 A2 A3 B1 B2 B3 P S Q [Li et al.(2020b)] X X X X X X X X [Lim et al.(2020a)] X X X X X X X X [Kairouz et al.(2021)] X X X X X X X X X X [Wahab et al.(2021)] X X X X X X X X X [Rahman et al.(2021)] X X X X X X X X [Bellavista et al.(2021)] X X X X X X X X [Gao et al.(2022)] X X X X X X X X X X [Ye et al.(2023)] X X X X X [Gecer and Garbinato(2024)] X X X X X X X X Ours
@@
Numerous surveys of federated learning in general or particular facets have been made available. Still, there is a need to offer a sensible and thorough taxonomy of the current state of energy-efficient (green) federated learning research difficulties and state-of-the-art, which is a significant and developing area of research for green-aware AI. Thus, we highlight the distinctive contributions of our work in Table 1 and evaluate the important contributions and limitations of previous relevant studies. The difficulties posed by federated learning are examined by Li et al.[Li et al.(2020b)] from four angles: statistical heterogeneity, system heterogeneity, communication efficiency, and privacy concerns. They also provide a quick overview of a number of potential future study areas. In the study of federated learning in mobile edge networks, Lim et al.[Lim et al.(2020a)] separated the approaches now in use into two categories: those that address the core issues with federated learning and those that apply federated learning to address edge computing issues. In addition to discussing current developments in federated learning, Kairouz et al.[Kairouz et al.(2021)] offered an overview of unresolved issues and concerns, such as effective communication, privacy protection, attack defense, and federated fairness. Wahab et al.[Wahab et al.(2021)] offer a detailed categorization system of current problems and solutions. The work mainly considered client selection and communication efficiency approaches. Rahaman et al.[Rahman et al.(2021)] divided the survey’s categories into data partitioning, FL architectures, algorithms/aggregation techniques, and personalization techniques as the implementation details and FL applications in various industries and domains, and communication cost, statistical heterogeneity, systems heterogeneity, and privacy/security as the main challenges. In [Bellavista et al.(2021)], the authors included information on the characteristics of federated environments, a unique taxonomy of decentralized learning methodologies, and a thorough explanation of the most pertinent and particular system-level contributions made by the surveyed solutions for non-IIDness, privacy, device heterogeneity, and poisoning defense. Data space, statistics, systems, and model heterogeneity in FL are covered by Gao et al.[Gao et al.(2022)], who also offer a taxonomy and an overview of the scenarios, objectives, and techniques associated with each heterogeneity challenge. In addition to offering multiple viewpoints to show the potential for future development of heterogeneous federated learning, the authors of in [Ye et al.(2023)] developed a new taxonomy to classify current heterogeneous federated learning approaches into three levels: data-level, model-level, and server-level. Gecer et al.[Gecer and Garbinato(2024)] conducted a survey of FL concerning efficiency, efficacy, privacy, robustness, and fairness. In order to compare and characterize federated learning solutions, the authors identified three abstract roles to present their conceptual model: aggregator, learner, and data creator. Some of the previous survey works [Ma et al.(2023), Pfeiffer et al.(2023), Qi et al.(2024)] are related to specific concerns such as non-IID data, heterogeneity, model aggregation etc.
@@
The primary research question (RQ1) focuses on popularly used FL variants and the roadmap of green FL. The RQ1 is addressed by sections Challenges and solutions in Federated Learning and Variants of Federated Learning Frameworks. The second research question (RQ2) is to offer a thorough and organized summary of all literature pertaining to green FL. Additionally, (RQ2) intends to demonstrate how adopting FL might be advantageous for the IoT application. The third research question (RQ3) is also driven by the need to address issues with energy efficiency and optimization in green FL IoT contexts. The RQ2 and RQ3 are addressed by Energy and Carbon Footprint Study in Federated Learning. The fourth research question (RQ4) offers final recommendations for a researcher in the green FL area that are mostly focused on issues with IoT applications. RQ4 is addressed by the sections FL based Green- Aware IoT Applications and Discussions & Future Scope. In this survey work, we explore FL approaches and the problems of an energy-efficient optimized distributed model, as well as challenges and opportunities with the application of the Internet of Things (IoT). It is pertinent to mention that due to ample evidence of security and privacy-preserving challenges and solutions, in this survey, we are not considering the security and privacy-preserving issues.
Main Contributions
@@
The research methodology used for this study is the “preferred reporting item for systematic reviews and meta-analyses” (PRISMA) [Page et al.(2021)]. The PRISMA technique is a widely used standard for providing evidence in systematic reviews, and it has been embraced by journals and organizations relevant to health. PRISMA methodologies provide several benefits, including enhancing the quality of the review, allowing the reader to evaluate its merits and shortcomings, simulating review procedures, and structuring and formatting the review using PRISMA headings. However, conducting a comprehensive examination and releasing it can take some time. It can also become rapidly obsolete, so it needs to be updated regularly to include all new primary data released since the start of the project.
@@
Challenges and solutions in Federated Learning
@@
1) System heterogeneity: System parameters like hardware, network connectivity, and energy limitations frequently differ between different clients [Wei et al.(2022)]. The straggler effect may result from this, namely in cross-device FL.
2) Statistical heterogeneity is related to client non-i.i.d. data [Zhu et al.(2021)]. For instance, different hospitals could have non-identifiable disease data because their patient population varies in geography and demographics. Different banks may have heterogeneous data because clients from various areas and nations may have varied financial backgrounds and investing styles. Three key approaches have been created by current studies to address statistical heterogeneity, and they are presented as follows:
@@
Personalization: Due to the drift of heterogeneous data, the global model generated by the server may hurt the performance and generalization of the local models on the clients. To solve this problem, personalization techniques modify the local models according to the local tasks. Various techniques are available for personalization. The approaches to this kind of personalization are either fine-tuning a global model on a local client distribution or pruning a global model. Nevertheless, these current approaches either predetermine network layers for fine-tuning, resulting in inadequate global information storage inside client models, or personalize at the expense of critical global knowledge retention [Tamirisa et al.(2023)]. Another approach for personalization is personalization layers. The negative consequences of statistical heterogeneity can be mitigated using personalization layers not subject to the federated averaging (FedAvg) process [Arivazhagan et al.(2019a)]. Another approach for personalization is knowledge distillation, a technique for knowledge transfer, to compare the results of a shared dataset and learn from each other’s inferred knowledge [Jeong and Kountouris(2023)].
Clustering: Though most FL frameworks operate under the assumption that there is only one global model, learning clients’ knowledge may not be adequate in extremely heterogeneous data sets. Clustering techniques address this problem by dividing the clients into several clusters and ensuring that each cluster upholds a server-client architecture. Specifically, similar clients are assigned to the same cluster based on the loss values or submitted model weights that indicate how similar they are [Ouyang et al.(2021)].
Optimization of Efficiency
@@
The literature mentions two main approaches to reduce communication overhead in federated learning. They are model compression and communication rounds [Ángel Morell et al.(2022)]. Client selection is another approach to reduce communication overhead as the involvement of more client participation makes convergent the global model earlier. Reducing the number of global iterations by synchronizing the maximum number of clients eventually lowers communication expenses and energy usage. [Thakur et al.(2023a)].
@@
Model compression can be achieved either using quantization or sparsification. Sparsification significantly lowers communication overhead, but it must be used carefully since transmitting too little information might negatively impact accuracy, while transmitting too much information can result in energy depletion and network congestion. Consequently, there is a trade-off between excessive and insufficient compression. It is costly to try to find the ideal balance through hyperparameter tweaking, and the tuning parameters that are produced are static and problem-specific. Furthermore, the majority of current compression techniques don’t take into account the varying communication costs associated with various technologies [Khirirat et al.(2021)]. Furthermore, because edge devices in the Internet of Things (IoT) have high energy constraints, the compression framework should be created to guarantee that the greatest amount of learning can occur while utilizing the least amount of energy to transmit the models—that is, to maximize the energy efficiency of edge nodes.
@@
Sparsification is a model compression technique for automatically determining the near-optimal communication and computation trade-off [Han et al.(2020)]. This trade-off is adjusted in the original FL method, federated averaging (FedAvg), by the number of local update rounds that occur in between each of the two communication (weight aggregation) rounds [McMahan et al.(2016)]. FedAvg either delivers all of the model parameters or nothing at all following each local model update step using gradient descent. Gradient sparsification (GS), a more balanced method that sends a sparse vector containing a selection of significant values from the complete gradient, has drawn interest recently in remote learning systems [Jiang and Agrawal(2018)]. In contrast to FedAvg’s ”send-all-or-nothing” strategy, GS offers more flexibility in managing the trade-off between computation and communication [Han et al.(2020)]. However, these approaches employ a single compression rate for every user and fail to take into account the communication heterogeneity in an actual FL system; consequently, these approaches are constrained by the worst communication capacity among users. Moreover, sparsification approaches lack adaptability and fail to leverage the redundant, similar information in users’ ML models for compression. In [Beitollahi et al.(2022)], the authors presented a novel approach to Dynamic Sparsification for Federated Learning (DSFL) that allows users to compress their local models according to their communication capacity at each iteration through the use of two novel methods for sparsification: layer-wise similarity sparsification (LSS) and extended top-K sparsification. DSFL is made possible by LSS to take advantage of the global redundant information present in users’ models through the sparsification process of Centralized Kernel Alignment (CKA) similarity. By allowing various values of the sparsification rate K for each user at each iteration, the extended top-K model sparsification approach enables DSFL to accommodate the diverse communication capability of user devices. Still, maintaining accuracy following a high ratio sparsification has never been addressed. To rectify the sparse gradient update process, a General Gradient Sparsification (GGS) framework is put forth in [Li et al.(2020a)] for adaptive optimizers. Gradient correction and batch normalization update with local gradients (BN-LG) are its two main techniques. The optimizer can correctly handle the accumulated unimportant gradients with gradient correction, improving the model’s convergence. Moreover, local gradient updates to the batch normalization layer can mitigate the effects of delayed gradients without adding to the communication overhead.
@@
Quantization is another widely used model compression technique. The FL setup presents several difficulties. For instance, depending on the requirements for their hardware design, various clients may have varying computational limitations. The hardware accelerator’s supported quantization bit-width is one realistically significant heterogeneous hardware attribute. A well-trained model ought to be quantized to a range of bit widths that are representative of the heterogeneous device landscape without exhibiting appreciable performance reduction. The goal of quantization robustness is to do this by training a single model with the constraint of robustness to different quantization bit-widths while at inference time, without the need for retraining or fine-tuning [Gupta et al.(2023a)]. The research hasn’t given enough thought to the energy requirements and hardware-induced limits for on-device learning, despite the effectiveness of common Federated optimization techniques like Federated Averaging (FedAvg) in FL. Enabling trained models to be quantized to different bit-widths based on the energy requirements and heterogeneous hardware designs across the federation is specifically a requirement for on-device learning.
@@
Nevertheless, these approaches, such as sparsification and quantization are pre-defined. They fail to capture redundant, correlated information across FL iterations, user device data, and ML model parameters. Furthermore, the error-correcting potential of the FL process is not entirely utilized by these methods. In [Beitollahi and Lu(2022)], the authors presented the Federated Learning with Autoencoder Compression (FLAC) method, which compresses user devices’ models for uplink transmission by taking advantage of FL’s redundant information and error-correcting capabilities. In the Training State, FLAC trains an autoencoder on the server to encode and decode user models. Later, in the Compression State, the autoencoder is sent to user devices to compress local models for further iterations. FLAC dynamically manages the autoencoder error by alternating between the Training State and Compression State to modify its autoencoder and its compression rate based on the error tolerance of the FL system, ensuring the convergence of the FL.
@@
Client Selection is another method to address the communication efficiency challenge in FL. Transferring a large number of updated model parameters over the uplink communication channel from the users to the server is a challenging task in FL which increases the communication complexity is a challenging task. Reducing the number of participating clients by scheduling policies [Amiri et al.(2020), Yang et al.(2020)] can help address this difficulty. In the literature, only a subset of clients can be scheduled for updates at each iteration due to bandwidth limitations. However, the random selection of a partial number of clients can reduce the convergence speed in the presence of heterogeneous clients. Moreover, selecting more clients ensures fast convergence in the presence of heterogeneous clients [Huang et al.(2023a)]. Hence, it is necessary to select participating clients to maintain the tradeoff between accuracy, convergence speed, and communication efficiency.
Computation Overhead
@@
Preliminaries and Problem Definition
@@
Variants of Federated Learning Frameworks